Express in formality the concentration of 28.4 per cent in NH3 by weight having a density of 0.808 grams per ml.


1
Expert's answer
2021-02-23T04:29:19-0500

CM=n/V

n=m/M

M (NH3) = 18 g/mol

Lets assume that we have 1 L of NH3 solution.

density = m/V

m = density x V

m (NH3)solution = 0.808 x 1 = 0.808 kg

%= (100 x m(solute))/(m(solute)+m(solvent))

m(NH3) = (28.4 x 808)/100 = 229.5 g

n(NH3) = 229.5/18 = 12.8 mol

CM(NH3) = 12.8/1 = 12.8 mol/L